A culture of 10 bacteria of Escherichia coli has a generation time of 30 minutes at 35 degrees Celsius. After 5 hours at this temperature, how many bacteria would you expect to see?"

a. "10,240"

b. 50

c. "1,000"

d. 320

e. "10,000,000,000"
Answer #1

Bacteria grow by binary fission, which means the number of bacteria doubles with each generation.

The initial number of bacteria 20 = 10

Generation time = 30 minutes, which means the bacterial cells double every 30 minutes.

We are asked to determine the number of bacteria after 5 hours.

Since the generation time is given as 30 minutes, we can calculate the number of generations of the bacterial culture in 5 hours or 5 * 60 = 300 minutes.

Number of generations = 300 / 30 = 10

Now, if there is a single bacterial cell, the number of bacteria after 10 generations would be 210 = 1024 bacterial cells.

Now, we would get the number of bacteria after 10 generations by multipying the initial number of bacteria with that of the number of generations: 10 * 210 = 10 * 1024 = 10,240.

Therefore, the right answer here would be option a.